What is the magnitude of the frictional force (in N) exerted on a plush toy of
0.125 kg which, when thrown on the floor with a speed of 3.60 m / s, travels 2.20 m
before stopping?
1
Expert's answer
2020-09-21T06:30:04-0400

Given

Mass of toy is "m=0.125 kg"

Initial velocity of toy is "v_0=3.60 m\/s"

Final velocity is "v=0"

Distance travelled is "d=2.20 m"

The acceleration of toy is

"a=\frac{v^2-v_0^2}{2d}=\frac{0^2-(3.60 m/s)^2}{2(2.20 m)}=-2.945 m/s^2"

According to Newton's second law,

"-f=ma"

So the frictional force is

"f=(0.125 kg)(2.945 m\/s^2)=0.37 N"